# Rolling with the Punches: Resilient Contrastive Pre-training under Non-Stationary Drift

### v2

@@ 3.2 Taming OOD Drift and Domain Drift in Downstream Tasks
Texture iNaturalist ImageNet-O OpenImage-O Overall
Methods PT ID AUROC FPR AUROC FPR AUROC FPR AUROC FPR AUROC FPR
MSP (hendrycks2016baseline) 71.3 77.1 90.7 43.7 60.8 90.6 84.3 61.8 76.8 68.3
Energy (liu2020energy) 54.1 86.3 76.6 72.7 61.6 81.0 71.1 74.0 65.9 78.5
MaxLog (hendrycks2019scaling) 67.2 78.0 89.9 45.6 61.7 88.6 82.7 62.5 75.4 68.7
KL (hendrycks2019scaling) 82.6 67.3 87.6 69.7 66.6 88.2 84.3 74.2 80.3 74.8
Residual (wang2022vim) 82.4 64.6 73.7 86.0 68.4 87.5 74.9 78.0 74.9 79.0
React (sun2021react) 62.1 80.5 91.2 38.7 63.7 81.0 80.4 60.4 74.3 65.2
Mahalanobis (lee2018simple) 84.9 66.1 84.9 81.6 71.5 88.9 84.2 74.7 81.4 77.8
ViM (wang2022vim) 83.5 62.7 77.8 81.7 71.0 86.6 78.3 74.6 77.7 76.4
MOODv1 (li2023rethinking) 93.0 30.9 98.8 5.9 86.8 63.2 95.5 26.5 93.5 31.6
MOODv2 (liMOODv2MaskedImage2024) -21K -1K 94.3 24.7 99.6 1.8 91.5 40.8 97.4 13.6 95.7 20.2
BYOL (grillBootstrapYourOwn2020) 58.4 92.1 80.1 77.5 52.0 92.2 65.3 81.3 64.0 85.8
DINO (caronEmergingPropertiesSelfSupervised2021) 72.9 80.7 85.7 65.3 74.2 80.2 74.7 83.0 76.9 77.3
DINO v2 (zhouDINOWMWorldModels2024) 67.1 85.1 78.9 88.4 63.1 88.3 73.9 85.4 70.7 86.8
MoCo v3 (chenEmpiricalStudyTraining2021) -LT -LT 81.5 63.9 80.1 68.6 65.0 91.0 69.9 84.2 74.1 76.9
Ours 83.1 57.8 88.0 67.7 70.1 82.9 75.5 78.3 79.2 71.7
Beyond that, to validate the generalization capability of our resilient contrastive pre-training under concept drift scenarios, we conducted experiments from perspectives of domain shift and out-of-distribution (OOD) detection.
In terms of domain shift generalization as exhibited in Table 3, our primary focus is to assess the extent of the drift environment on the feature representation constructed during pre-training. Therefore, we select three subsets of ImageNet (russakovskyImagenetLargeScale2015) to validate the domain shift performance, namely, ImageNet-V2 (rechtImageNetClassifiersGeneralize2019), ImageNet-Sketch (wangLearningRobustGlobal2019) and ImageNet-R (hendrycksManyFacesRobustness2021). Our results achieve superior results across three datasets, investigating the effectiveness of the proposed resilient contrastive pre-training method. It corroborates that, the feature space we construct in pre-training abstracts and captures the essential information of images, while disregarding any unexpected interference from the concept drift within the data stream. Moreover, judging from the results of Few split in ImageNet-Sketch and ImageNet-R, the majority of contrastive pre-training methods struggle to differentiate the tail categories. This difficulty arises from two main challenges. First, the bias induced by tail drift tends to skew models toward favoring head categories. On the flip side, the small sample size of tail categories limits the ability of pre-training methods to accurately extract their fundamental features, a problem that is exacerbated in the presence of domain shifts. Accordingly, our outstanding results demonstrate our proficiency in mitigating tail drift and extracting crucial features from limited samples.
Concerning out-of-distribution (OOD) detection, we also evaluate the capability of our resilient contrastive pre-training approach to delineate the boundaries of the feature space. While exhibiting slightly lower performance than DINO on ImageNet-O, we significantly outperform contrastive pre-training methods on the other three out-of-distribution datasets, and surpasses other approaches in overall performance. It demonstrates that our model possesses the benefits of intra-class compactness and inter-class separability in feature representation. Meanwhile, it is worth noting that in larger-scale balanced pre-training datasets, such as ImageNet-21k (ridnikImageNet21KPretrainingMasses2021), our approach outperforms numerous traditional OOD detection methods in overall performance, such as ViM (wang2022vim). It underscores our resilient contrastive pre-training methods can effectively alleviate concept drift bias in the model while also enhancing the characterization of the model’s decision boundary.
@@ 3.3 Boosting Pre-training Feature Embedding
Table 5: Evaluation results of different contrastive learning strategies in the stage of pre-training, from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories. The cosine metric is utilized to measure these distances, which is expressed as average degrees. We compare our methods with other contrastive pre-training models, including BYOL, DINO, DINO v2 and MoCo v3. ImageNet-LT is utilized for pre-training. The best-performing models are highlighted in red.
Pre-training Many Medium Few All
ID Intra-class Compactness $\downarrow$
BYOL (grillBootstrapYourOwn2020) 53.1 59.1 59.9 56.9
DINO (caronEmergingPropertiesSelfSupervised2021) 34.2 51.4 59.0 45.9
DINO v2 (zhouDINOWMWorldModels2024) 31.2 50.1 57.8 43.9
MoCo v3 (chenEmpiricalStudyTraining2021) 29.1 47.8 57.8 42.8
Ours 28.5 46.5 55.8 40.9
ID Inter-class Separability $\uparrow$
BYOL (grillBootstrapYourOwn2020) 81.5 79.2 78.3 80.0
DINO (caronEmergingPropertiesSelfSupervised2021) 89.0 88.4 87.6 88.5
DINO v2 (zhouDINOWMWorldModels2024) 88.9 88.1 87.2 88.3
MoCo v3 (chenEmpiricalStudyTraining2021) 89.3 88.8 88.2 88.9
Ours 89.4 89.0 88.4 89.1
ID vs. OOD Separability $\uparrow$
BYOL (grillBootstrapYourOwn2020) 75.7 71.9 70.3 73.1
DINO (caronEmergingPropertiesSelfSupervised2021) 86.2 84.4 81.4 84.7
DINO v2 (zhouDINOWMWorldModels2024) 86.0 83.7 80.5 84.1
MoCo v3 (chenEmpiricalStudyTraining2021) 86.9 85.7 83.0 85.8
Ours 87.1 86.1 83.4 86.1
The intra-class compactness results demonstrate that our model effectively validates the efficacy of the proposed resilient contrastive pre-training in delineating feature boundaries, especially in long-tailed scenarios. Moreover, the inter-class separability among different splits is very similar, suggesting that the pre-training primarily influences intra-class compactness rather than inter-class separability. Furthermore, with the exception of BYOL, the inter-class separability of the other contrastive pre-training methods is nearly the same. This observation underscores the significance of negative samples in expanding inter-class separation. It also demonstrates that the bias is induced by accumulated drift within momentum updates, where features of head categories dominated the whole model under tailed drift. In the context of ID vs. OOD separability, we attained the optimal results, signifying that our feature space exhibits distinct feature boundaries. In addition, we find that the performance of ID vs. OOD separability aligns with the ID intra-class compactness, suggesting that enhancing the model’s performance hinges on mitigating data stream drift to enhance feature extraction capabilities.
@@ 3.4 Scaling with Drift Pre-training
Table 6: Evaluation results of scaling ability of our resilient contrastive pre-training on ImageNet-LT with ViT-Small/16, ViT-Base/16 and ViT-Large/16. Top-1 accuracy is utilized to evaluate the performance.
Backbones ImageNet-LT
Many Medium Few All
ViT-S/16 67.7 39.6 11.6 46.4
ViT-B/16 72.2 43.1 16.0 50.3
ViT-L/16 73.8 45.0 16.9 52.0
To showcase the scalability of our resilient contrastive pre-training approach in scenarios involving concept drift, we evaluate the efficiency and scalability of our model in Table 6. As the number of parameters in ViT increases, our model exhibits a scalable effect, leading to performance improvements with larger models in downstream long-tailed classification tasks. In comparison to ViT-S/16, our proposed resilient contrastive learning method achieves nearly a 4% enhancement with ViT-B/16. This suggests that we can effectively train larger models to attain superior performance in the face of concept drift within data streams.
More critically, our experiments reveal that pre-trained models exhibit consistent scaling behavior even on datasets with great distribution drift. It demonstrates that our proposed resilient contrastive pre-training can effectively leverage larger-scale, minimally curated datasets, thereby substantially expanding the pool of usable data for pre-training without relying on complicated cleaning procedures.
@@ 4 Conclusion and Outlooks
In this paper, we present resilient contrastive pre-training, a novel, straightforward and effective pre-training paradigm tailored for concept drift data stream. We employ causal inference theory to methodically examine the source of bias in the momentum update of contrastive pre-training and put forward a causal interventional objective to mitigate this bias within the drifting data stream. By virtue of this objective, resilient contrastive pre-training is devised to counteract the unpredictable distribution changes occurring within the data stream.
We hope that our work will inspire future advancements in contrastive pre-training paradigm, specifically addressing biases originating from real-world data challenges. In future research, we will leverage causal inference to analyze mixure of concept drift in pre-training.
Impact Statement
This paper presents work whose goal is to advance the field of machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related Works
A.1 Concept Drift
@@
In the comprehensive survey conducted by Lu et al. (luLearningConceptDrift2019; lu2020data), existing approaches for concept drift handling are systematically classified into three primary categories: error rate-based methods (wangSelfadaptiveEnsembleUser2024; jiaoDynamicEnsembleSelection2024), data distribution-based methods (yang2024adaptingmultimodallargelanguage; cerqueiraSTUDDStudentTeacher2023), and multiple hypothesis-based techniques (yu2024online; yuLearnadaptConceptDrift2022). Our proposed methodology falls under the category of distribution-based concept drift adaptation approaches. These distribution-driven techniques distinguish themselves by not only enabling precise drift identification through explicit statistical distribution analysis but also providing a multidimensional characterization of drift patterns - including temporal occurrence detection, affected feature space localization, and quantitative severity assessment. This dual capability of detection coupled with comprehensive drift diagnostics establishes distribution-based methods as particularly valuable for developing adaptive systems that require interpretable drift understanding and targeted model adjustments.
@@
Besides, Online Boosting Adaptive Learning (OBAL) method (yu2024online) is proposed to address the challenges of concept drift and negative transfer in multistream classification. OBAL employs a dual-phase approach: an initial model is built using the Adaptive Covariate Shift Adaptation (AdaCOSA) algorithm to handle covariate shifts and learn dynamic correlations among streams. In the online phase, a Gaussian Mixture Model-based weighting mechanism is integrated to manage asynchronous drift. Meanwhile, CDMLLM (yang2024adaptingmultimodallargelanguage) reveals that vision-language models suffer significant bias from concept drift during pre-training and fine-tuning. To address this, the authors propose a unified concept drift framework integrating T-distribution-based adaptation for long-tailed calibration and explicit OOD detection, demonstrating enhanced robustness in open-world multi-modal alignment through systematic distribution modelling. Beyond on drift detection in single data stream, GDDM (yuDetectingGroupConcept2023) focuses on addressing group concept drift across multiple data streams, where individual drifts may go undetected due to subtle changes in underlying distributions. The proposed method introduces a distribution-free test statistic to detect concept drift in these complex scenarios. By designing an online learning algorithm for streaming data, the approach accurately identifies concept drift caused by hypothesis testing. Beyond that, DDG-DA (liDDGDataDistributionGeneration2022) proactively models predictable factors influencing environmental evolution. The approach involves training a predictor to estimate future data distribution, using this information to generate training samples, and then training models on the generated data. By leveraging predictable factors to forecast data distribution changes, DDG-DA aims to enhance model performance in handling concept drift in streaming data. Furthermore, STUDD(cerqueiraSTUDDStudentTeacher2023) proposes a teacher-student paradigm to enable unsupervised drift detection through deviation analysis of their predictive consistency. This approach leverages model disagreement as a proxy signal, bypassing dependency on ground-truth labels during deployment while maintaining detection sensitivity.
A.2 Causal Inference
@@
Recently, increasing researchers have incorporated causal inference into deep-learning models, especially in large models. Deconfounded Image Captioning (DIC) (yangDeconfoundedImageCaptioning2023) is proposed to address dataset bias in vision-language models through a causal lens, that integrates backdoor and front-door adjustments for systematic bias mitigation. The framework provides principled causal analysis of spurious correlations in multimodal alignment, offering theoretical grounding for decomposing bias sources through structured interventions. Likewise, aiming for spurious correlations induced by visual and linguistic biases during training, CIIC (liuShowDeconfoundTell2022) is proposed as a causal intervention framework combining an Interventional Object Detector (IOD) and Interventional Transformer Decoder (ITD) guided by structural causal models. By applying backdoor adjustment through IOD’s feature disentanglement and ITD’s dual de-confounding mechanism, their approach systematically mitigates confounding effects across encoding and decoding stages, demonstrating enhanced generalization through causal correlation modeling. Similarly, targeting multi-hop fact verification bias in the large language model, Causal Walk (zhangCausalWalkDebiasing2024) is proposed, a front-door adjustment framework that disentangles complex spurious correlations in evidence chains. The method models reasoning paths as mediators in structural causal models, decomposing causal effects via random walk-based treatment-mediator estimation and geometric mean-based mediator-outcome approximation. By integrating adversarial and symmetric datasets synthesized with large language models, the approach demonstrates superior debiasing performance.
@@
Additionally, causal inference is widely used in representation learning. Comprehensive Interventional Distillation (CID) (dengComprehensiveKnowledgeDistillation2021) integrates causal intervention with class-aware representation alignment. By reinterpreting teacher logits as contextual confounders and applying counterfactual pruning through structural causal models, CID systematically disentangles beneficial semantic patterns from dataset-specific biases. This approach demonstrates enhanced generalization through bias-invariant knowledge transfer. Besides, De-confound-TDE (tangLongTailedClassificationKeeping2021) establishes a causal framework for long-tailed classification, identifying SGD momentum as a paradoxical confounder that simultaneously harms tail-class predictions while benefiting representation learning. Through causal intervention during training and counterfactual reasoning at inference, the method disentangles momentum’s detrimental bias from its beneficial mediation effects. Meanwhile, CCM (miaoDomainGeneralizationContrastive2022) is proposed to address domain generalization through causal invariance principles. The framework integrates front-door adjustment with contrastive learning to quantify stable causal effects across domains, explicitly modeling domain shifts via a three-stage process: domain-conditioned supervision for feature correlation, causal effect measurement through structured path manipulation, and contrastive clustering for class-consistent representations. Similarly, CIRL (lvCausalityInspiredRepresentation2022) advances domain generalization through causal factorization, proposing a structural causal model that decomposes inputs into invariant causal mechanisms and domain-specific non-causal factors. It enforces three critical properties: causal/non-causal separation, joint independence, and causal sufficiency for classification.
@@
Besides, C2L (choiC2LCausallyContrastive2022) addresses model fragility to spurious patterns through contrastive counterfactual synthesis, proposing a collective decision framework that aggregates predictions across generated counterfactual sets. Unlike conventional augmentation limited by dataset-inherent biases, this approach probabilistically supervises causal invariance through distributional consensus, demonstrating enhanced robustness against attribution bias and domain shifts. Furthermore, ABCD (rohekarCausalInterpretationSelfAttention2023) establishes a causal interpretation of Transformer self-attention mechanisms, modeling them as structural equation estimators that capture conditional independence relations through partial correlation analysis in deep attention layers. This framework enables zero-shot causal discovery over input sequences while accounting for latent confounders, effectively repurposing pre-trained models for causal graph inference. What’s more, Causal Attention (CATT) (yangCausalAttentionVisionLanguage2021) implements front-door adjustment to address confounding bias in attention mechanisms via dual-path processing of In-Sample and Cross-Sample Attention. By forcibly integrating external sample contexts through CS-ATT while maintaining standard attention conventions, CATT dynamically mitigates spurious correlations without requiring confounder specification.
A.3 Contrastive Pre-training

### v3

@@ 3.3 Boosting Pre-training Feature Embedding
The intra-class compactness results demonstrate that our model effectively validates the efficacy of the proposed RCP in delineating feature boundaries, especially in long-tailed scenarios. Moreover, the inter-class separability among different splits is very similar, suggesting that the pre-training primarily influences intra-class compactness rather than inter-class separability. Furthermore, with the exception of BYOL, the inter-class separability of the other contrastive pre-training methods is nearly the same. This observation underscores the significance of negative samples in expanding inter-class separation. It also demonstrates that the bias is induced by accumulated drift within momentum updates, where features of head categories dominated the whole model under tailed drift. In the context of ID vs. OOD separability, we attained the optimal results, signifying that our feature space exhibits distinct feature boundaries. In addition, we find that the performance of ID vs. OOD separability aligns with the ID intra-class compactness, suggesting that enhancing the model’s performance hinges on mitigating data stream drift to enhance feature extraction capabilities.
@@ 3.4 Scaling with Drift Pre-training
Table 6: Evaluation results of scaling ability of our RCP on ImageNet-LT with various ViT models. Top-1 accuracy is utilized to evaluate the performance.
Backbones Many Medium Few All
ViT-Small/16 67.7 39.6 11.6 46.4
ViT-Base/16 72.2 43.1 16.0 50.3
ViT-Large/16 73.8 45.0 16.9 52.0
To present the scalability of our RCP approach in scenarios involving concept drift, we evaluate the efficiency and scalability of our model in Table 6. As the number of parameters increases, our model exhibits a scalable effect, leading to performance improvements with larger models in downstream classification tasks. In comparison to ViT-S/16, our proposed method achieves nearly a 4% enhancement with ViT-B/16. This suggests that we can effectively train larger models to attain superior performance in the face of concept drift within data streams.
More critically, our experiments reveal that pre-trained models exhibit consistent scaling behavior even on datasets with great distribution drift. It demonstrates that our proposed RCP can effectively leverage larger-scale, minimally curated datasets, thereby substantially expanding the pool of usable data for pre-training without relying on complicated cleaning procedures.
3.5 Ablation on Adaptation Window Length
Table 7: Ablation experiments on adaptation window length within the linear probing task on ImageNet-LT. We also provide the results of MoCo v3 for comparison.
Window Many Medium Few All
MoCo v3 [chenEmpiricalStudyTraining2021] 4,096 45.8 18.7 0.0 27.1
Ours 2,048 42.1 12.8 1.1 22.4
Ours 4,096 47.9 20.1 3.2 28.5
Ours 9,600 52.7 24.0 7.9 32.8
Furthermore, we also conduct ablation experiments on adaptation window length, as exhibited in Table 7. It demonstrates that our approach relies on a wider drift adaptation window, with superior results under a larger one. In particular, our method consistently outperforms MoCo v3 on Few Split, demonstrating the robustness of our method against non-stationary drift.
@@ 4 Conclusion and Outlooks
In this paper, we present resilient contrastive pre-training (RCP), a novel, straightforward and effective pre-training paradigm tailored for concept drift data stream. We employ causal inference theory to methodically examine the source of bias in the momentum update of contrastive pre-training and put forward a causal interventional objective to mitigate this bias within the drifting data stream. By virtue of this objective, RCP is devised to counteract the unpredictable distribution changes occurring within the data stream.
[…]
Supplementary Material
5 Related Works
5.1 Concept Drift
@@
In the comprehensive survey conducted by Lu et al. [luLearningConceptDrift2019, lu2020data], existing approaches for concept drift handling are systematically classified into three primary categories: error rate-based methods [wangSelfadaptiveEnsembleUser2024, jiaoDynamicEnsembleSelection2024], data distribution-based methods [yang2024adaptingmultimodallargelanguage, cerqueiraSTUDDStudentTeacher2023], and multiple hypothesis-based techniques [yu2024online, yuLearnadaptConceptDrift2022, yu2025drift]. Our proposed methodology falls under the category of distribution-based concept drift adaptation approaches. These distribution-driven techniques distinguish themselves by not only enabling precise drift identification through explicit statistical distribution analysis but also providing a multidimensional characterization of drift patterns - including temporal occurrence detection, affected feature space localization, and quantitative severity assessment. This dual capability of detection coupled with comprehensive drift diagnostics establishes distribution-based methods as particularly valuable for developing adaptive systems that require interpretable drift understanding and targeted model adjustments.
@@
Besides, Online Boosting Adaptive Learning (OBAL) method [yu2024online] is proposed to address the challenges of concept drift and negative transfer in multistream classification. OBAL employs a dual-phase approach: an initial model is built using the Adaptive Covariate Shift Adaptation (AdaCOSA) algorithm to handle covariate shifts and learn dynamic correlations among streams. In the online phase, a Gaussian Mixture Model-based weighting mechanism is integrated to manage asynchronous drift. Meanwhile, CDMLLM [yang2024adaptingmultimodallargelanguage] reveals that vision-language models suffer significant bias from concept drift during pre-training and fine-tuning. To address this, the authors propose a unified concept drift framework integrating T-distribution-based adaptation for long-tailed calibration and explicit OOD detection, demonstrating enhanced robustness in open-world multi-modal alignment through systematic distribution modelling. Beyond on drift detection in single data stream, GDDM [yuDetectingGroupConcept2023] focuses on addressing group concept drift across multiple data streams, where individual drifts may go undetected due to subtle changes in underlying distributions. The proposed method introduces a distribution-free test statistic to detect concept drift in these complex scenarios. By designing an online learning algorithm for streaming data, the approach accurately identifies concept drift caused by hypothesis testing. Beyond that, DDG-DA [liDDGDataDistributionGeneration2022] proactively models predictable factors influencing environmental evolution. The approach involves training a predictor to estimate future data distribution, using this information to generate training samples, and then training models on the generated data. By leveraging predictable factors to forecast data distribution changes, DDG-DA aims to enhance model performance in handling concept drift in streaming data. Furthermore, STUDD[cerqueiraSTUDDStudentTeacher2023] proposes a teacher-student paradigm to enable unsupervised drift detection through deviation analysis of their predictive consistency. This approach leverages model disagreement as a proxy signal, bypassing dependency on ground-truth labels during deployment while maintaining detection sensitivity.
5.2 Causal Inference
@@
Recently, increasing researchers have incorporated causal inference into deep-learning models, especially in large models. Deconfounded Image Captioning (DIC) [yangDeconfoundedImageCaptioning2023] is proposed to address dataset bias in vision-language models through a causal lens, that integrates backdoor and front-door adjustments for systematic bias mitigation. The framework provides principled causal analysis of spurious correlations in multimodal alignment, offering theoretical grounding for decomposing bias sources through structured interventions. Likewise, aiming for spurious correlations induced by visual and linguistic biases during training, CIIC [liuShowDeconfoundTell2022] is proposed as a causal intervention framework combining an Interventional Object Detector (IOD) and Interventional Transformer Decoder (ITD) guided by structural causal models. By applying backdoor adjustment through IOD’s feature disentanglement and ITD’s dual de-confounding mechanism, their approach systematically mitigates confounding effects across encoding and decoding stages, demonstrating enhanced generalization through causal correlation modeling. Similarly, targeting multi-hop fact verification bias in the large language model, Causal Walk [zhangCausalWalkDebiasing2024] is proposed, a front-door adjustment framework that disentangles complex spurious correlations in evidence chains. The method models reasoning paths as mediators in structural causal models, decomposing causal effects via random walk-based treatment-mediator estimation and geometric mean-based mediator-outcome approximation. By integrating adversarial and symmetric datasets synthesized with large language models, the approach demonstrates superior debiasing performance.
@@
Additionally, causal inference is widely used in representation learning. Comprehensive Interventional Distillation (CID) [dengComprehensiveKnowledgeDistillation2021] integrates causal intervention with class-aware representation alignment. By reinterpreting teacher logits as contextual confounders and applying counterfactual pruning through structural causal models, CID systematically disentangles beneficial semantic patterns from dataset-specific biases. This approach demonstrates enhanced generalization through bias-invariant knowledge transfer. Besides, De-confound-TDE [tangLongTailedClassificationKeeping2021] establishes a causal framework for long-tailed classification, identifying SGD momentum as a paradoxical confounder that simultaneously harms tail-class predictions while benefiting representation learning. Through causal intervention during training and counterfactual reasoning at inference, the method disentangles momentum’s detrimental bias from its beneficial mediation effects. Meanwhile, CCM [miaoDomainGeneralizationContrastive2022] is proposed to address domain generalization through causal invariance principles. The framework integrates front-door adjustment with contrastive learning to quantify stable causal effects across domains, explicitly modeling domain shifts via a three-stage process: domain-conditioned supervision for feature correlation, causal effect measurement through structured path manipulation, and contrastive clustering for class-consistent representations. Similarly, CIRL [lvCausalityInspiredRepresentation2022] advances domain generalization through causal factorization, proposing a structural causal model that decomposes inputs into invariant causal mechanisms and domain-specific non-causal factors. It enforces three critical properties: causal/non-causal separation, joint independence, and causal sufficiency for classification.
@@
Besides, C2L [choiC2LCausallyContrastive2022] addresses model fragility to spurious patterns through contrastive counterfactual synthesis, proposing a collective decision framework that aggregates predictions across generated counterfactual sets. Unlike conventional augmentation limited by dataset-inherent biases, this approach probabilistically supervises causal invariance through distributional consensus, demonstrating enhanced robustness against attribution bias and domain shifts. Furthermore, ABCD [rohekarCausalInterpretationSelfAttention2023] establishes a causal interpretation of Transformer self-attention mechanisms, modeling them as structural equation estimators that capture conditional independence relations through partial correlation analysis in deep attention layers. This framework enables zero-shot causal discovery over input sequences while accounting for latent confounders, effectively repurposing pre-trained models for causal graph inference. What’s more, Causal Attention (CATT) [yangCausalAttentionVisionLanguage2021] implements front-door adjustment to address confounding bias in attention mechanisms via dual-path processing of In-Sample and Cross-Sample Attention. By forcibly integrating external sample contexts through CS-ATT while maintaining standard attention conventions, CATT dynamically mitigates spurious correlations without requiring confounder specification.
5.3 Contrastive Pre-training